Calculate the maximum energy and the speed


of the emitted electrons when sodium is


illuminated by a radiation of wavelength


150nm. What is the least frequency of radiation


(threshold frequency) for which electrons are


emitted?


(Take work function of sodium as 2.0eV).


Given that me= 9.109 x 10-31kg,


mn= 1.675 x 10-27kg, and


mp= 1.672 x 10-27kg

1
Expert's answer
2022-01-24T08:44:50-0500

"h\frac c\lambda=A+E,"

"E=h\frac c\lambda-A=6.3~eV,"

"v=\sqrt{\frac {2E}m}=1.49\cdot 10^6~\frac ms ."

"\nu_{min}=\frac Ah=482.7~GHz."